2 HW Problem Show Me How business, $40,000 3 Purchased supplies on account, $3,200 4 Paid creditor on account $1,750 $420 8 Determined that the cost of supplies used was $1,400 9 Withdrew cash for
6 Paid automobile expenses (ncluding rental charge) for month, $1,880, and $420 7 Paid office salaries, $5,000 8 Determined that the cost of supplies used was $1,400 Required: 1.Journalize entries for transactions Jan. 1 through 9. Refer to the Chart of Accounts for exact wording of account titles 2. Post the journal entries to the T accounts, selecting the appropriate date to the left of each amount to identify the transactions. Determine the account balances after all posting is complete. Accounts containing only a single entry do not need a balance 3. Prepare an 4 Determine the following d trial balance as of January 31, 2019 a Amount of total revenue recorded in the ledger b. Amount of total expenses recorded in the iedger c Amount of net income for January 5. Determine the increase or decrease in gwners equity for January eck My Work
Answer #1

Requirement 1: Journal entries

Jan, 2019

1 . Cash ac Dr 40000

To Capital ac 40000

(Being capital introduced into business by Sharon)

2. Rent on office building ac Dr . 6000

To Cash ac 6000

(Being rent on office equipment paid)

3. Supplies on hand ac Dr. 3200

To Accounts payable 3200

(Being supplies purchased on account)

4. Accounts payable . Dr . 1750

To Cash. 1750

( Being accounts payable settled in cash to the extent of $1750)

5. Cash ac Dr 18250

To Earned fees ac 18250

( Being fees earned and received received in cash)

6. Automobile expense ac. Dr. 1880

Miscellaneous expense ac. Dr. 420

To Cash ac 2300

(Being expenses paid)

7. Salaries ac. Dr. 5000

To Cash ac. 5000

( Being salary paid in cash)

8. Profit and loss ac. Dr 1400

To Supplies on hand ac. 1400

(Being cost of supplies ascertained and charged to profit and loss account assumed to be passed on 8th itself)

9. Capital ac. Dr. 2000

To Cash ac. 2000

( Being cash withdrawn by owner)

Note: profit and loss account means income statement

Requirement 2: Ledger accounts

Cash Account

Jan 2019 $ Jan 2019 $
1 To capital 40000 2 By rent on office equipment 6000
5 To Earned fees 18250 4 By accounts payable 1750
6 By automobile expenses 1880
6 By miscellaneous expenses 420
7 By salary 5000
9 By capital 2000
31 By balance c/d 41200
58250 58250

Capital account

Jan 2019 $ Jan 2019 $
9 to cash 2000 1 by cash 40000
31 to balance c/d 38000
40000 40000

Rent on office building

Jan 2019 $ jan 2019 $
To Cash 6000

Supplies on hand account

Jan 2019 $ jan 2019 $
3 to accounts payable 3200 8 by profit and loss account 1400
By balance cd 1800
3200 3200

Accounts payable account

Jab 2019 $ Jan 2019 $
4 to cash 1750 3 by supplies 3200
31 to balance cd 1450
3200 3200

Earned fees account

Jan 2019 $ Jan 2019 $
5 by cash 18250

Automobile expense account

Jan 2019 $ Jan 2019 $
6 To Cash 1880

Miscellaneous expenses account

Jan 2019 $ Jan 2019 $
To Cash 420

Salary account

Jan 2019 $ Jan 2019 $
To Cash 5000

Requirement 3: unadjusted trial balance

Dr Cr
Cash 41200
Capital 38000
Supplies 3200
Accounts payable 1450
Rent on office equipment 6000
Earned fees 18250
Automobile expenses 1880
Miscellaneous expenses 420
Salary 5000
Total 57700 57700

Requirement 4:

A. Total revenue recorded in ledger= earned fees= $18250

B . Total expense recorded in ledger= 6000+ 1400+ 1880+ 420+ 5000= $14700

C. Net income= A-B= $3550

Note the cost of supplies is taken to calculate the expenses. Closing balance of supplies does not apply to this period.

Requirement 5: there is only a withdrawal of $2000. Therefore owners equity has a balance of $38000 now.